OpenStudy (anonymous):

. Simplify. 8t + 6r − 3t + 2r ? show work

OpenStudy (saifoo.khan):

5t + 8t

OpenStudy (anonymous):

8-3=5 6+2=8

OpenStudy (anonymous):

show work

OpenStudy (saifoo.khan):

thts a single step problem..

OpenStudy (anonymous):

8t-3t+6r+2r =5t+8r

OpenStudy (anonymous):

saifoo your first answer was wrong >:(